# Training Software Engineering Agents and Verifiers with SWE-Gym

### v1

@@ 4.1.2 Verifier for Agent Scaffold with Specialized Workflow
Figure 6: Scaling inference-time compute for MoatlessTools Agents with learned verifiers. We set temperature $t=0.5$ during sampling.
@@
For MoatlessTools Agents with specialized workflows, given that it doesn’t have a turn-taking action-observation trajectory like OpenHands CodeActAgent, we prepare verifier inputs through a parsing process. This process combines task descriptions, relevant agent context, and generated patches, training the model to output a single confidence token indicating task success.. We adapt the context extractor from Zhang et al. (2024a) and provide the prompt template in §B.4.
@@
We train 7B and 32B verifiers using on-policy trajectories from the last round of sampling, applying LoRA (Hu et al., 2022) for regularization. To address data bias, we cap positive data points per instance at 2 and balance the dataset by subsampling failure cases to match successful ones.
@@
Result. We evaluate the verifiers by sampling the policy 8 times at temperature 0.5. As shown in Fig. 6, these verifiers enable effective scaling across model sizes - the 7B agent-verifier system improves from 10% to 13.3% success rate on SWE-Bench Lite, while the 32B system improves from 19.7% to 26.3%. The 7B verifier plateaus after N=4 samples when ranking trajectories from both 7B and 32B agents. In contrast, the 32B verifier continues improving even at N=8, suggesting model size affects scaling behavior.
@@ 4.2 Training-Time Scaling with Data
In this subsection, we study the impact of training data scaling on agent performance. We identify three types of data scaling for sampled trajectories: (1) Random Scaling (No Dedup.), which simply increases the total number of training trajectories regardless of uniqueness (Fig. 7); (2) Unique Instance Scaling (Dedup.), which scales the number of distinct instances by randomly selecting one success trajectory per instance (Fig. 8); and (3) Repository Scaling (Dedup.), which samples trajectories by sequentially including all instances from one repository at a time, sorted by repository name, to study the impact of repository-level diversity.
@@
Setup. Using OpenHands (Wang et al., 2024c), we evaluate these scaling approaches on SWE-Bench Verified: random scaling on the full trajectory dataset from §3.2 (491 trajectories), unique instance scaling on these trajectories deduplicated by instance ID (294 trajectories), and repository-based scaling where we sort repositories alphabetically and include all trajectories from each repository in order (e.g., first 25% contains complete trajectories from the first N repositories). We compare models trained on 25%, 50%, and 100% of the corresponding dataset for each approach. Please refer to Tab. 7 for detailed statistics of these datasets.
@@
Model performance scales well with more training examples. As illustrated in Fig. 7, we observe favorable scaling behavior: there is consistent improvements in model resolve rate as training data increases, particularly for the 32B model. These results suggest that SWE-Gym’s current size and repository diversity are likely not performance bottlenecks - further improvements could likely be achieved by allocating more computing resources to sampling more training trajectories.
Figure 7: Model performance scaling with training data size. The x-axis shows the percentage of training data used in log base 2 scale.
@@
Similar scaling trends suggest instance and repository diversity is not yet a bottleneck. Fig. 7 reveals comparable overall performance between different scaling approaches up to where deduplication takes effect. While Random Scaling (No Dedup.) achieves higher final performance, this is likely due to having more trajectories (491 vs 294) rather than better scaling efficiency. Among deduplicated approaches, Repository Scaling shows stronger initial performance at 25% data, suggesting that complete repository coverage may provide more coherent learning signals early in training. These results suggest that the repository and instance diversity of SWE-Gym is not yet a bottleneck - further improvements could likely be achieved by simply sampling more agent trajectory data for traning, regardless of duplication or repository distribution.
Figure 8: Comparison of three data sampling approaches: without deduplication, repository-based sampling, and random sampling (§4.2). All variants use the 32B model evaluated on SWE-Bench Verified.
5 Related Work
5.1 Agents that Solve Github Issues
In this paper, we focus on software engineering agents designed to automatically resolve GitHub issues, specifically in the context of SWE-Bench (Jimenez et al., 2024). These agents operate in a well-defined setting: given a GitHub issue and its associated code repository, they generate a valid code modification (i.e., git diff patch) to address the issue. The correctness of these modifications is verified through the execution of a test suite written by human developers.
Existing agent designs can be categorized by the number of human priors injected into the agent workflow: (1) specialized workflow: Xia et al. (2024); Örwall (2024); Zhang et al. (2024b); Chen et al. (2024) build specialized workflow with human-defined stages (e.g., localization, code edit, patch re-ranking selection), where an LM is iteratively prompted to produce the final results. While restricting the LM to work on pre-defined subtasks, this approach effectively reduces the task horizon and alleviates the need for long-term planning. However, these specialized workflows face limitations: they require significant human engineering to define appropriate stages and transitions, may not generalize well to novel issue types outside their designed workflow, and can be brittle when intermediate steps fail. (2) general-purpose prompting: Systems such as Yang et al. (2024); Wang et al. (2024c) rely on general prompting approaches, including methods like ReAct (Yao et al., 2023) and CodeAct (Wang et al., 2024b). These systems primarily depend on the LM itself to plan through long horizons and produce the next action based on an evolving history of actions and observations, without enforcing a fixed workflow or relying heavily on human pre-defined stages. While more flexible, these general approaches place higher demands on the underlying LM’s capabilities and can be computationally expensive due to multiple rounds of interaction with a growing interaction history.
Most existing successful methods are based on prompting proprietary models and crafting specialized workflow to workaround existing limitations of those models. This starkly contrasts the success of learning-based approaches in other domains (Silver et al., 2017; Akkaya et al., 2019), where a system learns from prior interactions and rewards to develop task competency progressively. We argue that proper training environments and baselines can help accelerate research in this direction. We verify that we can use SWE-Gym to build strong agents through learning.
5.2 Environments for Training Software Agents
There is no existing dataset suitable for training software engineering agents. Existing datasets often exhibit distinct constraints. SWE-Bench (Jimenez et al., 2024) is widely used for evaluating software engineering performance. The training split lacks executable environments and success signals present in the evaluation split, making its usefulness limited for model training. HumanEval (Chen et al., 2021) is designed for standalone code generation tasks, akin to coding competitions. Therefore, it falls short of addressing the complex challenges inherent in real-world, repository-level software engineering tasks, which involve thousands of files, millions of lines of code, and tasks such as bug fixing, feature development, and system optimization. Similarly, R2E (Jain et al., 2024) is a small evaluation dataset with 246 instances and, due to its synthetic nature, lacks the realism and complexity in real world software engineering scenario. There is also limited empirical evidence if this environment is useful for training. Our proposed SWE-Gym instead uses real-world GitHub issue bodies as task instructions, and uses the associated unit tests for validation. This approach introduces a realistic and complex task formulations, aligning closely with real-world software engineering challenges.
5.3 Post-training: From Chatbots, Reasoners to Agents
Post-training, which finetunes a pre-trained language model, usually through supervised or reinforcement learning, has proven highly effective in improving model performance across various domains. RLHF (Ouyang et al., 2022) has become a standard method for adapting language models into chatbots, improving both performance and alignment of the model (Team, 2024). In mathematical reasoning, two standard datasets, MATH (Hendrycks et al., 2021b) and GSM-8K (Cobbe et al., 2021), have question and answer pairs for both training and evaluation, which enables researchers to explore methods to train both policy, and verifiers (reward models) (Cobbe et al., 2021; Wang et al., 2024a).
Earlier works  (Wang et al., 2024b; Chen et al., 2023; Zeng et al., 2023; Wu et al., 2024) shows distilling agent trajectories from stronger models consistently improves weaker models. Recent works shift towards self-improving methods. Xi et al. (2024); Zhai et al. (2024); Bai et al. (2024) show how RL or rejection sampling fine-tuning, guided by a reward signal, can enable models to self-improve without reliance on a more capable teacher model.
Despite the wide success, post-training generally rely on either expert demonstration data or a training environment with reliable reward signals, which is largely missing in the software engineering domain. This limitation has contributed to a reliance on prompting-based methods with proprietary language models, as reflected in most top-performing approaches on SWE-bench (Jimenez et al., 2024). Our work addresses this gap with SWE-Gym, a training environment grounded in real-world software engineering tasks, using associated expert written tests as the reward signal. Our experiments verify that we can use SWE-Gym to build strong SWE agents without any prompt engineering. Concurrent with our work, Ma et al. (2024) and Golubev et al. (2024) both study the training of software engineering agents, with the latter also exploring verifiers training and dataset construction. We discuss the key differences between these concurrent works in §A.
6 Conclusions
@@
In this paper, we introduce SWE-Gym, the first open training environment that bridges critical gaps in enabling scalable learning for software engineering agents. By combining real-world Python tasks with repository-level context, pre-configured execution environments, and test verifications, SWE-Gym will be a foundation for advancing LM agent training research. Through extensive experiments, we demonstrate that SWE-Gym enables both agent and verifier models to achieve significant improvements in resolving complex software tasks. Our findings highlight the scalability of these approaches, revealing potential for continuous performance gains with increased compute.
7 Acknowledgments
@@
We thank John Yang and Ofir Press for helpful discussions, and John Yang for assistance in reproducing data analysis results from SWE-Bench. We thank Modal Labs [^6] https://modal.com/ for the GPU compute support through its Academic Credits Program.
@@ Appendix B Experiment Details
B.2 Training Details.
OpenHands Agent Experiment. We use torchtune (PyTorch Team, 2024) for full parameter fine-tuning with a learning rate of 1e-4, maximum 5 epochs, global batch size of 8, max context length of 32768. We fine-tuned both 7B, 14B, and 32B variant of the model, and experiments were performed with 2-8x NVIDIA H100 80G GPU on modal (Modal, 2024). The only exception is in the main experiment of §4.1.1, where we use LoRA (Hu et al., 2022) (29.8% @$8$) via Unsloth library (Unsloth Team, 2024) to train the verifier for max 2 epochs, while other hyper-parameter stays the same.
MoatlessTools Agent Experiment. All MoatlessTools models are trained with a context window of 10240. For experiments with the 7B model, we use torchtune to train the policy model with full-finetuning using 4 H100 GPUs. We set batch size to 8, learning rate to $2\times 10^{-5}$, and train for 5 epochs.
@@
B.3 Details of OpenHands Trajectory Sampling

### v2

@@ 5.1.2 Verifier for Specialized Workflow
Figure 4: Scaling inference-time compute for MoatlessTools Agents (32B) with learned verifiers on SWE-Bench Lite. Temperature $t=0.5$.
@@
For MoatlessTools agents with specialized workflows, given that it doesn’t have a turn-taking action-observation trajectory like OpenHands CodeActAgent, we prepare verifier inputs through a parsing process adopted from Zhang et al. (2024a), which combines task descriptions, relevant agent context, and generated patches. [^9] We provide the prompt template in §B.5. We train the verifier to map from this input to a single token indicating task success.
@@
Following the training procedure described in §5.1.1, we train 7B and 32B verifiers using on-policy trajectories from the last (2nd round of sampling, applying LoRA (Hu et al., 2022). To address the easy-data bias in the training dataset, we cap the number of successful trajectories per instance at two and balance the data by subsampling failure cases to match the same number of successful ones.
@@
Results. We evaluate the verifiers by sampling from an agent policy with $k=$ 8 at temperature 0.5. As shown in Fig. 4 and Fig. 7, these verifiers enable effective scaling across verifier and policy sizes: the 7B verifier improves from 10 to 13.3% resolution rate on SWE-Bench Lite when paired with a 7B policy, while the 32B verifier improves from 19.7 to 26.3% when paired with a 32B policy. The 7B verifier plateaus after $k=$ 4 samples when ranking trajectories from both 7B and 32B agents. In contrast, the 32B verifier continues improving even at $k=$ 8, suggesting that verifier size significantly affects scaling behavior.
@@ 5.2 Training-Time Scaling with Data
We then examine how scaling the amount of training data affects agent performance using 491 sampled trajectories from §4.2. We simulate three scaling methods through subsampling: (1) Scaling trajectories, where trajectories are randomly dropped (Fig. 5); (2) Scaling unique task instances, where only one successful trajectory per task instance is selected (Fig. 9); and (3) Scaling repositories, which sequentially includes all instances from each repository to assess repository-level diversity.
@@
Setup. Using OpenHands (Wang et al., 2024c) and the fine-tuning approach described in §4.2, we evaluate these scaling approaches on SWE-Bench Verified: scaling the number of trajectories, by subsampling from the full trajectory dataset from §4.2 (at most 491 trajectories); unique instance scaling on these trajectories deduplicated by instance ID (at most 294 trajectories), and repository-based scaling where we sort repositories alphabetically and include all trajectories from each repository in order (e.g., first 25% contains complete trajectories from the first N repositories). We compare models trained on 25%, 50%, and 100% of the full dataset for each approach, sampling training subsets using the methods described above for each scaling approach. [^10] Tab. 7 contains detailed statistics of these datasets.
@@
Scaling trends suggest instance and repository diversity is not yet a bottleneck. Fig. 5 demonstrates substantial scaling behavior, with consistent improvements in resolution rate as the number of training trajectories randomly increases, particularly for the 32B model. These results suggest that SWE-Gym’s current size and repository diversity are likely not performance bottlenecks - further improvements could likely be achieved by allocating more computing resources to sampling more training trajectories.
Figure 5: Scaling effects of increasing the number of randomly sampled trajectories for training.
@@
Fig. 9 reveals comparable overall performance between different scaling approaches up to where deduplication takes effect. While Random Scaling (No Dedup.) achieves higher final performance, this is likely due to having more trajectories (491 vs 294) rather than better scaling efficiency. Among deduplicated approaches, Repository Scaling shows stronger initial performance at 25% data, suggesting that complete repository coverage may provide more coherent learning signals early in training. These results suggest that the repository and instance diversity of SWE-Gym is not yet a bottleneck - further improvements could likely be achieved by simply sampling more agent trajectory data for traning, regardless of duplication or repository distribution.
6 Conclusions, Limitations, and Future Work
@@
In this paper, we introduce SWE-Gym, the first training environment that addresses critical gaps in enabling scalable learning for software engineering agents. By combining real-world Python tasks with repository-level context, pre-configured execution environments, and test verifications, SWE-Gym will be a foundation for advancing LM agent training research. Through extensive experiments, we demonstrate that SWE-Gym enables both agent and verifier models to achieve significant improvements in resolving complex software tasks. Our findings highlight the scalability of these approaches, revealing potential for continuous performance gains with increased compute.
We see many research directions that we are excited to explore in the future:
  1. 1.
    Automatic Environment Synthesis SWE-Gym, while effective, is limited by its environment diversity, including the number of repositories, types of tasks, and programming languages. We view environment synthesis—via automated environment creation, test-case generation, or task generation—as a critical next step.
  2. 2.
    Self-Improvement with Reinforcement Learning Despite notable progress, our self-improvement results are modest. Training language model agents with large-scale online reinforcement learning is a promising direction for further improvements.
  3. 3.
    Human-Agent Interaction Current SWE settings focus solely on task completion, neglecting human-in-the-loop collaboration, which is essential for real-world software engineering. Methods like user simulation or learning from offline human-agent interaction data might offer ways for developing collaborative agents that align with human.
Impact Statement
This work presents SWE-Gym, an environment for training software engineering agents, with strong empirical results on its effectiveness. We discuss a few important societal implications to consider. First, improving automated software engineering capabilities could increase developer’s productivity and accessibility across industries. Although current models are primarily research artifacts and not yet production-ready, they can support critical open-source infrastructure and potentially make software development more accessible. Secondly, as these agents become more capable, they may impact software engineering jobs and require careful consideration around code ownership, licensing, and attribution. Additionally, while we focus on legitimate software engineering tasks, similar techniques could potentially be misused to automate the creation of malicious code. We encourage future work to further explore frameworks for responsible deployment of software engineering agents, including considerations around security, safety, and economic impacts.
Acknowledgments
@@
We thank John Yang and Ofir Press for helpful discussions, and John Yang for assistance in reproducing data analysis results from SWE-Bench. We thank Modal Labs [^11] https://modal.com/ for the GPU compute support through its Academic Credits Program. XW and HJ are partially supported by U.S. DARPA ITM Program No. FA8650-23-C-7316. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix A Comparison with Concurrent Works
[…]
0 (Zero-shot) 0 56.3 7.0
1 36 37.3 9.0
2 62 29.0 9.7
3 82 43.7 7.7
No Cap (All) 172 30.7 9.3
Table 6: resolution rate and empty patch rate on SWE-Bench Lite with a 7B model trained using different instance capping strategies (Cap).
Figure 7: Scaling inference-time compute for MoatlessTools Agents (7B and 32B) with their corresponding learned verifiers. Temperature $t=0.5$.
Figure 8: Ablation study for verifier training (§5.1.1). Performances are evaluated on SWE-Bench Verified. Both the agent and the verifier are Qwen2.5-Coder-Instruct-32B model fine-tuned on the corresponding dataset. OpenHands (Wang et al., 2024c) is used as the agent scaffold.
Figure 9: Comparison of three data sampling approaches using 32B LM: scaling trajectories (dedup.), scaling unique task instances, and scaling repositories (§5.2).
@@ Appendix B Experiment Details
B.2 OpenHands Agent Experiments
During training, we use OpenHands’s remote runtime (Neubig & Wang, 2024) feature to execute agent trajectories in parallel on SWE-Gym. We use torchtune (PyTorch Team, 2024) for full parameter fine-tuning with a learning rate of 1e-4, maximum 5 epochs, global batch size of 8, max context length of 32768. We fine-tuned both 7B, 14B, and 32B variant of the model, and experiments were performed with 2-8x NVIDIA H100 80G GPU on modal (Modal, 2024). The only exception is in the main experiment of §5.1.1, where we use LoRA (Hu et al., 2022) (29.8% @$8$) via Unsloth library (Unsloth Team, 2024) to train the verifier for max 2 epochs, while other hyper-parameter stays the same.
Inference during evaluation is bounded by either 100 interaction turns or the base LM’s 32k context window length, whichever is reached first.
B.3 MoatlessTools Agent Experiments
All MoatlessTools models are trained with a context window of 10240. For experiments with the 7B model, we use torchtune to train the policy model with full-finetuning using 4 H100 GPUs. We set batch size to 8, learning rate to $2\times 10^{-5}$, and train for 5 epochs.
@@ B.3 MoatlessTools Agent Experiments
B.4 Details of OpenHands Trajectory Sampling